# An Efficient Randomized Fixed-Precision Algorithm for Tensor Singular Value Decomposition

### 17 Jan 2024 v10

@@ IV Computational complexity
V Simulations
In this section, we test the proposed randomized algorithm using synthetic and real-world data. All numerical simulations were performed on a laptop computer with 2.60 GHz Intel(R) Core(TM) i7-5600U processor and 8GB memory. In all our simulations, we set the power iteration $q=1$. The compression ratio is defined as $\frac{\rm Number\,of\,parameters\,of\,the\,original\,tenors}{\rm Number\,of\,parameters\,in\,the\,compressed\,model}$. We have used some Matlab function from the toolbox https://github.com/canyilu/Tensor-tensor-product-toolbox.
@@ V Simulations
Then, we applied the truncated t-SVD with the mentioned tubal ranks to the underlying data tensors. The running times of the proposed algorithm and the t-SVD are reported in Table II. It is seen that the Algorithm 4 outperforms the truncated t-SVD in terms of running time.
@@
VI Conclusion and future works
@@
VII Acknowledgements
After acceptance of the paper, the author found similar incremental algorithms for computation of the t-SVD in [39, 40]. Thanks Dr. Ugochukwu Ugwu for bringing these works to the author’s attention. The author would like to thank Stanislav Abukhovich for his help in the implementation of the algorithms. He introduced the author the Matlab function cat to perform the tensor concatenation operation. He also pointed out constructive comments on the proof of Theorem 3. The author is also indebted to three anonymous reviewers for their constructive and useful comments which have greatly improved the quality of the paper. The author was partially supported by the Ministry of Education and Science of the Russian Federation (grant 075.10.2021.068).

### 08 Apr 2024 v11

@@ V Computational complexity
VI Simulations
In this section, we test the proposed randomized algorithm using synthetic and real-world data. All numerical simulations were performed on a laptop computer with 2.60 GHz Intel(R) Core(TM) i7-5600U processor and 8GB memory. In all our simulations, we set the power iteration $q=1$. The compression ratio is defined as $\frac{\rm Number\,of\,parameters\,of\,the\,original\,tenors}{\rm Number\,of\,parameters\,in\,the\,compressed\,model}$. We have used some Matlab function from the toolbox https://github.com/canyilu/Tensor-tensor-product-toolbox. The codes are available at https://github.com/SalmanAhmadi-Asl/Fixed-Precision-t-SVD.
@@ VI Simulations
Then, we applied the truncated t-SVD with the mentioned tubal ranks to the underlying data tensors. The running times of the proposed algorithm and the truncated t-SVD are reported in Table II. It is seen that the Algorithm 4 outperforms the truncated t-SVD in terms of running time.
@@
VII Conclusion and future works
@@
VIII Acknowledgements
After acceptance of the paper, the author found similar incremental algorithms for the computation of the t-SVD in [39, 40]. Thanks Dr. Ugochukwu Ugwu for bringing these works to the author’s attention. The author would like to thank Stanislav Abukhovich for his help in the implementation of the algorithms. He introduced the author the Matlab function cat to perform the tensor concatenation operation. He also pointed out constructive comments on the proof of Theorem 3. The author is also indebted to three anonymous reviewers for their constructive and useful comments which have greatly improved the quality of the paper. The author was partially supported by the Ministry of Education and Science of the Russian Federation (grant 075.10.2021.068).